Shift summary for owl 8/4/99

Jason Hamm (jhamm@physics.arizona.edu)
Wed, 4 Aug 1999 05:34:41 -0700 (MST)


Shift Summary for owl shift, 08/4/99

On Shift: J. Hamm, A. Tripathi

Hours of beam available at all intensities: 3.5, only 2.0 of which was
known to be on target
Hours of beam available at nominal intensity: 3.5 (2.0)
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): .75

Accounting of lost time: aftermath of beamline power glitch
========================================================================
Runs ended this shift:

14097 - Laser scan

Tapes brought to FCC this shift: none
========================================================================
Accesses during shift: none
========================================================================
Overview of shift: high-stress, no data
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
0:00 Still no beam after beamline power glitch

0:05 Continuing to verify HV MF's after power glitch.
MF 15 (regenerator) shows channels 8-11 volts from
database. Expert notified.
hv1440.tcl, lvmux_csi SDAQ processes died sometime
sometime during the glitch - these are restarted.

1:04 JRL finishes fruit debugging for the night.

1:45 Start laser scan 14097

2:27 End laser scan 14097

3:30 Channel 13 says "beam shortly" - in anticipation
we start beam_832 run 14100

3:35 Beam arrives, but NM2SEED1&2, NM2WC1 appear to be
dead --> no beam position information. We're
instructed to push the crash switch and keep the
beam stop in until we're sure the beam is on
target. Beam continues to come, but we and MCR
have no way of knowing if it's any good.Stop 14100

5:05 NM2SEED2, NM2WC1 finally appear and look on target.
Since these are the 2 inputs to autotune, we're
set to go. Start beam_832 14101.

We immediately suffer from ktevana/pipeline
errors. After various combinations of kfrends,
resets, and repeater cyclings, we finally get
triggers through. Online histos look OK, until...

5:45 Filling out the run summary sheet we notice that
NM2ABS2 is set to stop 0 (0") and reading back
stop 2 (10"), both of which are incorrect - it
should be at stop 1 (20"). Setting it to 1
does not help, epicure still reads back 2.Stop run
and page Rick C. He notifies ops and advises us
to put the beam stop in. Meanwhile we sit and
wait to find out where the absorber is really
sitting.

6:45 MCR calls and informs us that the absorber is now
tracking correctly - it is both set and reads back
at stop 1. The problem was apparently a dirty
readback card.

6:50 Start beam_832 run 14103. Start getting "scaler
l1raw_b1&4/sem out of limits" errors on every
spill. Reg/SEM high at ~1.5, DC's drawing a
lot of current (~90 uA for DC1). Still trying to
track down problem at end of shift.

7:15 Turns out that absorber was still stuck in some
intermediate setting, even though epicure gave us
a readback of 1. Rick F. moved it from stop 1 to
2 and back to 1 - all errors go away. Run 14103
stopped and is useless.

Things are NOW running smoothly...
========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift